OpenStudy (anonymous):

Solve for X: radical (3x+7) + 4= X+7

OpenStudy (anonymous):

(3x+7) + 4= X+7 3x+11 = x+7 3x-x = 7-11 do it now

OpenStudy (anonymous):

hmm how did you get 3x+11=x+7

OpenStudy (anonymous):

as x+7 is given in your question so its ok but (3x+7) + 4 removing the brackets we get 3x+7+4 now 7+4 = 11 thus 3x+11 = x+7

OpenStudy (mathstudent55):